) Following the uncertainty in UK economy, primarily as a result of the UK government’s failure to negotiate a trade deal with the European Union, Inger.com which operates in the pharmaceutical sector has experienced increasing price pressure. Despite this, Inger expects to see an above normal dividend growth rate of 20 per cent, but only for the last five years. Stock market analyst at RIBU Investment expects this above normal growth rate to continue for a further 5 years before levelling off at a normal rate of 6 per cent. Inger’s last dividend was €0.50 per share, and as a result of the less stable UK economic environment, analysts expect investors to demand a risk premium of 6 percent. It is assumed that the risk-free rate is 4 per cent. (i) Determine the current value of Inger’s stock on the basis that its required rate of return is 15 per cent and (ii) given the risk premium, will investors buy the stock?
Answer #1
Year Dividend
0 €      0.50
1 €      0.60
2 €      0.72
3 €      0.86
4 €      1.04
5 €      1.24
6 €      1.32
TV €   14.65
V0 $10.13

Forecast the future dividends given the growth rates.

D1 = D0 x (1 + g) = 0.50 x (1 + 20%) = 0.6 and so on.

Terminal Value (TV) is the price in year 5 = D5 x (1 + g) / (r - g) = 1.24 x (1 + 6%) / (15 % - 6%) = $14.65

Value today, V0 = D1 / (1 + r) + D2 / (1 + r)^2 + D3 / (1 + r)^3 + D4 / (1 + r)^4 + (D5 + TV) / (1 + r)^5

= 0.6 / 1.15 + 0.72 / 1.15^2 + 0.86 / 1.15^3 + 1.04 / 1.15^4 + (1.24 + 14.65) / 1.15^5

= $10.13